How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 31/2 yards long, and 5 pieces, each 21/4 yards long, are removed?
Answer by fractalier(6550)

This is found by subtracting...
24 - 3(3 1/2) - 5(2 1/4) =
24 - 10 1/2 - 11 1/4 =
24 - 21 3/4 =
2 1/4 yards are left.